OpenStudy (anonymous):

i dont know what this wants me to do solve the following equations, find all solutions (x-4)(x+5)=0 x^2-8x+12=0

OpenStudy (anonymous):

Solutions are x = 4 and x = -5

OpenStudy (anonymous):

for the first one

OpenStudy (akshay_budhkar):

Solve both the equations separately. the common answers is the solution to your problem

OpenStudy (anonymous):

thanks! wb the next one...?

OpenStudy (akshay_budhkar):

6 and 2

OpenStudy (anonymous):

how?

OpenStudy (akshay_budhkar):

x^2-6x-2x+12=0 x(x-6)-2(x-6) (x-2)(x-6)=0 therefore x=6 or x=2

jimthompson5910 (jim_thompson5910):

I'm going to factor the left side. To do that, I need to find 2 numbers that multiply to 12 but add to -8. These two numbers are -6 and -2 x^2-8x+12=0 x^2-6x-2x+12=0 x(x-6)-2(x-6)=0 (x-2)(x-6)=0 x-2=0 or x-6=0 x=2 or x=6

OpenStudy (anonymous):

To read it off you want it factorised which is how the first equation is represented. Then when one of the factors equals zero the whole equation will equal zero. That's why you can just read it off... for (x-4) if x = 4, 4-4 = 0 Therefore i know one solution is x= 4, for (x+5) if x = -5 then those brackets will equal zero -5 +5 = 0, therefore i know the other solution is x=-5 With the second equation you need to factorise it (get it in the form or (x+a)(x+b)) before you can read off the solutions. Hope that makes sense.

OpenStudy (anonymous):

thanks a million

OpenStudy (akshay_budhkar):

welcome a billion lol